# Calibrating Generative Models

### v1

@@ C.1 Precise Statement
Since the maximum entropy problem is not specific to generative model calibration, we present it in a more general setting. Our presentation builds on standard results from exponential families and convex analysis. We recommend Wainwright & Jordan (2008) for relevant background.
@@
Observe $\mathcal{A}$ is a bounded, linear map defined on this space. And $\psi$ and $g$ are convex. By Fenchel-Rockafellar duality (Borwein & Zhu, 2005, Theorem 4.4.2), weak duality holds for the maximum entropy problem and its dual (16).
Wainwright & Jordan (2008, Theorem 3.3) states that $\nabla_{\boldsymbol{\alpha}}A_{P}$ is a surjective mapping from $\Xi$ onto $\text{relint}(\mathcal{M})$. Hence, there exists $\boldsymbol{\alpha}^{\ast}\in\Xi$ for which $\nabla_{\boldsymbol{\alpha}}A_{P}(\boldsymbol{\alpha}^{\ast})=\boldsymbol{h}^{\ast}$. The value of the dual at $\boldsymbol{\alpha}^{\ast}$ is
@@ C.2 Estimating the Maximum Entropy Solution
If Section˜C.2 holds, then the exponential family (17) is said to be minimal. An exponential family for which Section˜C.1 holds is minimal if and only if the log-normalizer $A_{P}(\boldsymbol{\alpha})$ is strictly convex on $\Xi$ (Wainwright & Jordan, 2008, Proposition 3.1).
@@
However, in order for $\widehat{\boldsymbol{\alpha}}_{N}$ to be well-defined, the interior point condition and uniqueness of the dual optimum must hold for the maximum entropy problem with $P=\frac{1}{N}\sum_{n=1}^{N}\delta_{\boldsymbol{x}_{n}}$. For this problem, these two conditions are straightforward to verify: (i) $\boldsymbol{h}^{\ast}$ lies in the the relative interior of the convex hull of $\{\boldsymbol{h}(\boldsymbol{x}_{n})\}_{n=1}^{N}$ and (ii) the empirical covariance matrix of $\{\boldsymbol{h}(\boldsymbol{x}_{n})\}_{n=1}^{N}$ has full rank. For the example we provided at the beginning of the section, conditions (i) and (ii) are satisfied if and only if $\{\boldsymbol{h}(\boldsymbol{x}_{n})\}=\{0,1\}$ and $\boldsymbol{h}^{\ast}\in(0,1)$.
It is possible for Assumptions C.1 and C.2 to hold for $p_{\theta_{\text{base}}}$ but not for $\frac{1}{N}\sum_{n=1}^{N}\delta_{\boldsymbol{x}_{n}}$. For our example, if $\{\boldsymbol{h}(\boldsymbol{x}_{n})\}=\{0\}$ and $\boldsymbol{h}^{\ast}=0$ (or $\{\boldsymbol{h}(\boldsymbol{x}_{n})\}=\{1\}$ and $\boldsymbol{h}^{\ast}=1$), then the maximum entropy solution exists and is equal to $Q^{\ast}=\frac{1}{N}\sum_{n=1}^{N}\delta_{\boldsymbol{x}_{n}}$, but every vector $\boldsymbol{\alpha}\in\mathbb{R}$ is optimal for the dual problem (16). We demonstrate in Section˜C.4 that the probability of this event approaches zero as the number of samples $N$ approaches infinity. However, we observe (e.g., Figure˜2B) that when the base model $p_{\theta_{\text{base}}}$ lies far from the maximum entropy solution $p_{\boldsymbol{\alpha}^{\ast}}$, estimating $\boldsymbol{\alpha}^{\ast}$ with small variance requires many samples, and may even be computationally intractable.
@@ C.3 Connection Between the Relax and Reward Losses
In convex analysis (e.g., Boyd & Vandenberghe, 2004; Ben-Tal & Nemirovski, 2023), (19) is known as a penalty problem.
@@
defined on Euclidean space. We then show that $\boldsymbol{\alpha}_{\lambda}$ is the unique solution to the dual problem, and we use this solution to construct a solution to the primal problem. Interestingly, $\boldsymbol{\alpha}_{\lambda}$ is the unique solution to the dual problem even when there is redundancy among the constraints $\boldsymbol{h}$ (i.e. Section˜C.2 does not hold).
@@
As in the proof of Theorem˜C.3, $\mathcal{A}$ is a bounded, linear map defined on this space, and $\psi$ and $g$ are convex. By Fenchel-Rockafellar duality (Borwein & Zhu, 2005, Theorem 4.4.2), weak duality holds for the problem (19) and its dual (20).
@@
Next, we show that as the regularization parameter $\lambda\to 0$, then $Q_{\lambda}$ achieves the minimum possible Euclidean norm constraint violation i.e. Euclidean norm difference between $\mathbb{E}_{Q_{\lambda}}[\boldsymbol{h}]$ and $\boldsymbol{h}^{\ast}$. We also give a finite $\lambda$ bound on the constraint violation.
@@
$\lim_{\lambda\to\infty}\|\mathbb{E}_{Q_{\lambda}}[\boldsymbol{h}(\boldsymbol{x})]-\boldsymbol{h}^{\ast}\|=\inf_{\begin{subarray}{c}Q\in\mathcal{P}(P)\\ \textrm{D}_{\textrm{KL}}\left(Q\;\|\;P\right)<\infty\end{subarray}}\|\mathbb{E}_{Q}[\boldsymbol{h}(\boldsymbol{x})]-\boldsymbol{h}^{\ast}\|.$
@@
Fix $\epsilon>0$ such that the Euclidean ball of radius $\epsilon$ centered at $\boldsymbol{\alpha}^{\ast}$ is contained in $\Xi$. By Section˜C.2, $F_{0}$ is strictly concave, since then $\nabla_{\boldsymbol{\alpha}}^{2}A_{P}(\boldsymbol{\alpha})$ positive definite for every $\boldsymbol{\alpha}\in\Xi$. Hence, there exists a $\kappa$ such that for all $\|\boldsymbol{\alpha}-\boldsymbol{\alpha}^{\ast}\|=\epsilon$,
$F_{0}(\boldsymbol{\alpha})>\kappa>F_{0}(\boldsymbol{\alpha}^{\ast}).$
This is because the left-hand side of the above inequality attains its minimum on the compact set $\|\boldsymbol{\alpha}-\boldsymbol{\alpha}^{\ast}\|=\epsilon$ and (ii) by strict concavity this minimum must be strictly greater than the right-hand side. Moreover, by uniform convergence of $F_{\lambda}$ to $F_{0}$, there exists $\lambda_{\epsilon}>0$ such that for all $\lambda<\lambda_{\epsilon}$ and all $\|\boldsymbol{\alpha}-\boldsymbol{\alpha}^{\ast}\|=\epsilon$
$F_{\lambda}(\boldsymbol{\alpha})>\kappa>F_{\lambda}(\boldsymbol{\alpha}^{\ast}).$ (22)
@@ C.4 Consistency and Asymptotic Normality
In this section, we discuss the large sample behavior of the estimator $\widehat{\boldsymbol{\alpha}}_{N}$ for the parameters $\boldsymbol{\alpha}^{\ast}$ of the reward loss. Under Assumptions C.1, C.1, and C.2, we show that as $N\to\infty$ and $d$ remains fixed, then $\widehat{\boldsymbol{\alpha}}_{N}$ is close to $\boldsymbol{\alpha}^{\ast}$ with high probability. And under stronger conditions, we demonstrate that $\widehat{\boldsymbol{\alpha}}_{N}$ has a limiting normal distribution. The asymptotics of $\widehat{\boldsymbol{\alpha}}_{N}$ have previously been studied in the subject of empirical likelihood (Qin & Lawless, 1994; Kitamura & Stutzer, 1997; Owen, 2001).
We first aim to establish that $\widehat{\boldsymbol{\alpha}}_{N}$ is close to $\boldsymbol{\alpha}^{\ast}$ with high probability as $N\to\infty$ i.e. $\widehat{\boldsymbol{\alpha}}_{N}$ is consistent for $\boldsymbol{\alpha}^{\ast}$. Define the functions
@@
according to Section˜C.4. In other words, the estimator $\widehat{\boldsymbol{\alpha}}_{N}$ has greatest asymptotic variance when $h_{b}$ is close to either 0 or $1$. Notice that we can compute the asymptotic variance of $\widehat{\boldsymbol{\alpha}}_{N}$ directly (i.e. without using Section˜C.4) by applying the delta method to $\bar{\boldsymbol{y}}_{N}$ and the function $z\mapsto\log(\frac{\boldsymbol{h}^{\ast}(1-z)}{(1-\boldsymbol{h}^{\ast})z})$, in which case we obtain the same value.
@@
Remark. Under weaker assumptions (Assumptions C.1 and C.1), the proof of Section˜C.4 implies that the log-normalizer $A_{P}(\boldsymbol{\alpha})$ has derivatives of all orders on $\Xi$. Indeed, this is a consequence of equation˜28, which implies that for every $\boldsymbol{\alpha}$, there exists a neighborhood of $\boldsymbol{\alpha}$ contained in $\Xi$ on which the $k$th derivative of $\exp\{r_{\boldsymbol{\alpha}}(\boldsymbol{x})\}$ is uniformly $p_{\theta_{\text{base}}}$-dominated. This allows one to exchange differentiation and integration in the definition of $A_{P}(\boldsymbol{\alpha})$.
Appendix D Simulation Experiments Additional Details
In this section, we provide details for our experiments calibrating mixture proportions in a product of GMMs (Section˜3). First, in Section˜D.1 we give background on continuous-time diffusion models, including how we sample from $p_{\theta}$ and compute densities $p_{\theta}/p$ with respect to a dominating measure $p$. This enables us to employ CGM-relax and CGM-reward for calibrating a pre-trained diffusion model. In Section˜D.2, we describe how the base diffusion model can be initialized to generate exact samples from a GMM or product of GMMs. Lastly, in Section˜D.3 we provide details regarding our implementation of the CGM calibration algorithm, including optimizer, neural network architecture, and hyperparameters.
@@ D.1 Continuous-time Diffusion Models
Computing densities. In order to employ CGM-relax and CGM-reward, $p_{\theta}$ and $p_{\theta_{\text{base}}}$ must have densities with respect to one another, and it must be possible to compute these densities. Girsanov’s Theorem (Cameron & Martin, 1944; Girsanov, 1960) provides conditions that guarantee these densities to exist and an expression for computing them.
@@
For high-dimensional diffusion models (e.g. Genie2 in our Section˜4.1 experiments) memory constraints preclude the naive approach to computing equation˜34 by instantiating each term in memory and simultaneously back-propogating gradients through all terms at once. However, because the gradient is a sum across time, it can computed in chunks. In practice, we divide $\{0,\ldots,T\}$ into $\lceil T/\texttt{chunk\_size}\rceil$ blocks of approximately equal size, where chunk_size is the largest chunk size that can fit into memory.
D.2 Initializing the Base Diffusion Model to sample a Gaussian mixture

### v2

@@ C.1 Precise Statement
Since the maximum entropy problem is not specific to generative model calibration, we present it in a more general setting. Our presentation builds on standard results from exponential families and convex analysis. We recommend Wainwright and Jordan (2008) for relevant background.
@@
Observe $\mathcal{A}$ is a bounded, linear map defined on this space. And $\psi$ and $g$ are convex. By Fenchel-Rockafellar duality (Borwein and Zhu, 2005, Theorem 4.4.2), weak duality holds for the maximum entropy problem and its dual (15).
Wainwright and Jordan (2008, Theorem 3.3) states that $\nabla_{\boldsymbol{\alpha}}A_{P}$ is a surjective mapping from $\Xi$ onto $\text{relint}(\mathcal{M})$. Hence, there exists $\boldsymbol{\alpha}^{\ast}\in\Xi$ for which $\nabla_{\boldsymbol{\alpha}}A_{P}(\boldsymbol{\alpha}^{\ast})=\boldsymbol{h}^{\ast}$. The value of the dual at $\boldsymbol{\alpha}^{\ast}$ is
@@ C.2 Estimating the Maximum Entropy Solution
If Section˜C.2 holds, then the exponential family (16) is said to be minimal. An exponential family for which Section˜C.1 holds is minimal if and only if the log-normalizer $A_{P}(\boldsymbol{\alpha})$ is strictly convex on $\Xi$ (Wainwright and Jordan, 2008, Proposition 3.1). Once we have imposed Section˜C.2, Section˜C.1 is equivalent to $\boldsymbol{h}^{\ast}\in\text{int}(\mathcal{M})$, since $\text{relint}(\mathcal{M})=\text{int}(\mathcal{M})$ as $\mathcal{M}$ is full-dimensional.
@@
However, in order for $\widehat{\boldsymbol{\alpha}}_{N}$ to be well-defined, the interior point condition and uniqueness of the dual optimum must hold for the maximum entropy problem with $P=\frac{1}{N}\sum_{n=1}^{N}\delta_{\boldsymbol{x}_{n}}$. For this problem, these two conditions are straightforward to verify: (i) $\boldsymbol{h}^{\ast}$ lies in the interior of the convex hull of $\{\boldsymbol{h}(\boldsymbol{x}_{n})\}_{n=1}^{N}$ and (ii) the empirical covariance matrix of $\{\boldsymbol{h}(\boldsymbol{x}_{n})\}_{n=1}^{N}$ has full rank. For the example we provided at the beginning of the section, conditions (i) and (ii) are satisfied if and only if $\{\boldsymbol{h}(\boldsymbol{x}_{n})\}=\{0,1\}$ and $\boldsymbol{h}^{\ast}\in(0,1)$.
It is possible for Assumptions C.1 and C.2 to hold for $p_{\theta_{\text{base}}}$ but not for $\frac{1}{N}\sum_{n=1}^{N}\delta_{\boldsymbol{x}_{n}}$. For our example, if $\{\boldsymbol{h}(\boldsymbol{x}_{n})\}=\{0\}$ and $\boldsymbol{h}^{\ast}=0$ (or $\{\boldsymbol{h}(\boldsymbol{x}_{n})\}=\{1\}$ and $\boldsymbol{h}^{\ast}=1$), then the maximum entropy solution exists and is equal to $Q^{\ast}=\frac{1}{N}\sum_{n=1}^{N}\delta_{\boldsymbol{x}_{n}}$, but every vector $\boldsymbol{\alpha}\in\mathbb{R}$ is optimal for the dual problem (15). We demonstrate in Section˜C.4 that the probability of this event approaches zero as the number of samples $N$ approaches infinity. However, we observe (e.g., Figure˜3B) that when the base model $p_{\theta_{\text{base}}}$ lies far from the maximum entropy solution $p_{\boldsymbol{\alpha}^{\ast}}$, estimating $\boldsymbol{\alpha}^{\ast}$ with small variance requires many samples, and may even be computationally intractable.
@@ C.3 Connection Between the Relax and Reward Losses
In convex analysis (e.g., Hestenes, 1969; Powell, 1969; Boyd and Vandenberghe, 2004; Ben-Tal and Nemirovski, 2023), (18) is known as a penalty problem.
@@
defined on Euclidean space. We then show that $\boldsymbol{\alpha}_{\lambda}$ is the unique solution to the dual problem, and we use this solution to construct a solution to the primal problem. Interestingly, $\boldsymbol{\alpha}_{\lambda}$ is the unique solution to the dual problem even when there is redundancy among the constraints $\boldsymbol{h}$ (i.e., Section˜C.2 does not hold).
@@
As in the proof of Theorem˜C.3, $\mathcal{A}$ is a bounded, linear map defined on this space, and $\psi$ and $g$ are convex. By Fenchel-Rockafellar duality (Borwein and Zhu, 2005, Theorem 4.4.2), weak duality holds for the problem (18) and its dual (19).
@@
Next, we show that as the regularization parameter $\lambda\to 0$, then $Q_{\lambda}$ achieves the minimum possible Euclidean norm constraint violation i.e., Euclidean norm difference between $\mathbb{E}_{Q_{\lambda}}[\boldsymbol{h}]$ and $\boldsymbol{h}^{\ast}$. We also give a finite $\lambda$ bound on the constraint violation.
@@
$\lim_{\lambda\to 0}\|\mathbb{E}_{Q_{\lambda}}[\boldsymbol{h}(\boldsymbol{x})]-\boldsymbol{h}^{\ast}\|=\inf_{\begin{subarray}{c}Q\in\mathcal{P}(P)\\ \textrm{D}_{\textrm{KL}}\left(Q\;\|\;P\right)<\infty\end{subarray}}\|\mathbb{E}_{Q}[\boldsymbol{h}(\boldsymbol{x})]-\boldsymbol{h}^{\ast}\|.$
@@
Fix $\epsilon>0$ such that the Euclidean ball of radius $\epsilon$ centered at $\boldsymbol{\alpha}^{\ast}$ is contained in $\Xi$. By Section˜C.2 $\nabla_{\boldsymbol{\alpha}}^{2}A_{P}(\boldsymbol{\alpha})$ positive definite for every $\boldsymbol{\alpha}\in\Xi$, which implies $F_{0}$ is strictly concave. Hence, there exists a $\kappa$ such that for all $\|\boldsymbol{\alpha}-\boldsymbol{\alpha}^{\ast}\|=\epsilon$,
$F_{0}(\boldsymbol{\alpha})<\kappa<F_{0}(\boldsymbol{\alpha}^{\ast}).$
This is because the left-hand side of the above inequality attains its maximum on the compact set $\|\boldsymbol{\alpha}-\boldsymbol{\alpha}^{\ast}\|=\epsilon$ and (ii) by strict concavity this maximum must be strictly less than the right-hand side. Moreover, by uniform convergence of $F_{\lambda}$ to $F_{0}$, there exists $\lambda_{\epsilon}>0$ such that for all $\lambda<\lambda_{\epsilon}$ and all $\|\boldsymbol{\alpha}-\boldsymbol{\alpha}^{\ast}\|=\epsilon$
$F_{\lambda}(\boldsymbol{\alpha})<\kappa<F_{\lambda}(\boldsymbol{\alpha}^{\ast}).$ (21)
@@
The convergence rate $\|\mathbb{E}_{Q_{\lambda}}[\boldsymbol{h}(\boldsymbol{x})]-\boldsymbol{h}^{\ast}\|=\mathcal{O}(\lambda)$ is standard for penalty methods (see e.g., Hestenes, 1969). In our argument, we additionally show that up to first order,
$\|\mathbb{E}_{Q_{\lambda}}[\boldsymbol{h}(\boldsymbol{x})]-\boldsymbol{h}^{\ast}\|\leq c\lambda\{1+\|\left(\lambda\mathbb{I}+2\nabla_{\boldsymbol{\alpha}}^{2}A_{P}(\boldsymbol{\alpha}^{\ast})\right)^{-1}\|\},\quad c\geq 0.$
Hence, for finite $0<\lambda\ll 1$, we would expect $\|\mathbb{E}_{Q_{\lambda}}[\boldsymbol{h}(\boldsymbol{x})]-\boldsymbol{h}^{\ast}\|$ to be small whenever the Fisher information matrix of the exponential family $\exp\{r_{\boldsymbol{\alpha}}(\boldsymbol{x})-A_{P}(\boldsymbol{x})\}P(d\boldsymbol{x})$ at $\boldsymbol{\alpha}=\boldsymbol{\alpha}^{*}$ has a large minimum eigenvalue.
@@ C.4 Consistency and Asymptotic Normality
In this section, we discuss the large sample behavior of the estimator $\widehat{\boldsymbol{\alpha}}_{N}$ for the parameters $\boldsymbol{\alpha}^{\ast}$ of the reward loss. Under Assumptions C.1, C.1, and C.2, we show that as $N\to\infty$ and $d$ remains fixed, then $\widehat{\boldsymbol{\alpha}}_{N}$ is close to $\boldsymbol{\alpha}^{\ast}$ with high probability. And under stronger conditions, we demonstrate that $\widehat{\boldsymbol{\alpha}}_{N}$ has a limiting normal distribution. The asymptotics of $\widehat{\boldsymbol{\alpha}}_{N}$ have previously been studied in the subject of empirical likelihood (Qin and Lawless, 1994; Kitamura and Stutzer, 1997; Owen, 2001).
We first aim to establish that $\widehat{\boldsymbol{\alpha}}_{N}$ is close to $\boldsymbol{\alpha}^{\ast}$ with high probability as $N\to\infty$ i.e., $\widehat{\boldsymbol{\alpha}}_{N}$ is consistent for $\boldsymbol{\alpha}^{\ast}$. Define the functions
@@
according to Section˜C.4. In other words, the estimator $\widehat{\boldsymbol{\alpha}}_{N}$ has greatest asymptotic variance when $h_{b}$ is close to either 0 or $1$. Notice that we can compute the asymptotic variance of $\widehat{\boldsymbol{\alpha}}_{N}$ directly (i.e., without using Section˜C.4) by applying the delta method to $\bar{\boldsymbol{y}}_{N}$ and the function $z\mapsto\log(\frac{\boldsymbol{h}^{\ast}(1-z)}{(1-\boldsymbol{h}^{\ast})z})$, in which case we obtain the same value.
@@
Remark. Under weaker assumptions (Assumptions C.1 and C.1), the proof of Section˜C.4 implies that the log-normalizer $A_{P}(\boldsymbol{\alpha})$ has derivatives of all orders on $\Xi$. Indeed, this is a consequence of equation˜27, which implies that for every $\boldsymbol{\alpha}$, there exists a neighborhood of $\boldsymbol{\alpha}$ contained in $\Xi$ on which the $k$th derivative of $\exp\{r_{\boldsymbol{\alpha}}(\boldsymbol{x})\}$ is uniformly $p_{\theta_{\text{base}}}$-dominated. This allows one to exchange differentiation and integration in the definition of $A_{P}(\boldsymbol{\alpha})$.
Appendix D Calibrating Continuous-time Diffusion Models
In this section, we provide details on how CGM can be applied to calibrate continuous-time diffusion models. First, in Section˜D.1 we give background on continuous-time diffusion models, including how we sample from $p_{\theta}$ and compute densities $p_{\theta}/p$ with respect to a dominating measure $p$. This enables us to employ CGM-relax and CGM-reward for calibrating a pre-trained diffusion model. In Section˜D.2, we discuss the setting where the calibration function depends only on the final time of the diffusion process. In this case, the solution to the maximum entropy problem is a diffusion process, and we provide a mathematical characterization of its drift. Finally, in Section˜D.3 we describe how the base diffusion model can be initialized to produce exact samples from a GMM or product of GMMs. This allows us to initialize the base model exactly in our synthetic experiments.
@@ D.1 Continuous-time Diffusion Models
Computing densities. In order to employ CGM-relax and CGM-reward, $p_{\theta}$ and $p_{\theta_{\text{base}}}$ must have densities with respect to one another, and it must be possible to compute these densities. Girsanov’s Theorem (Cameron and Martin, 1944; Girsanov, 1960) provides conditions that guarantee these densities to exist and an expression for computing them.
@@
For high-dimensional diffusion models (e.g. Genie2 in our Section˜4.1 experiments) memory constraints preclude the naive approach to computing equation˜33 by instantiating each term in memory and simultaneously back-propogating gradients through all terms at once. However, because the gradient is a sum across time, it can computed in chunks. In practice, we divide $\{0,\ldots,T\}$ into $\lceil T/\texttt{chunk\_size}\rceil$ blocks of approximately equal size, where chunk_size is the largest chunk size that can fit into memory.
D.2 Solution to the Maximum Entropy Problem
When the base model $p_{\theta_{\text{base}}}(\boldsymbol{x})$ constitutes a continuous-time diffusion model (29) satisfying certain regularity properties, and the constraint function $\boldsymbol{h}$ depends only on the path at time $t=1$, there exists a closed-form solution to the maximum entropy problem (4).
Let $p$ be the law of an SDE having diffusion coefficient $\sigma$ and initial distribution $p_{\text{init}}^{\prime}(\boldsymbol{x}(0))$; this is necessary for $p\ll p_{\theta_{\text{base}}}$ by Girsanov’s Theorem (Theorem˜D.1). By the chain rule for the KL divergence, the objective for the maximum entropy problem defined on the full path measures is
$\textrm{D}_{\textrm{KL}}\left(p(\boldsymbol{x})\;\|\;p_{\theta_{\text{base}}}(\boldsymbol{x})\right)$
$=$ $\textrm{D}_{\textrm{KL}}\left(p(\boldsymbol{x}(1))\;\|\;p_{\theta_{\text{base}}}(\boldsymbol{x}(1))\right)+\mathbb{E}_{p(\boldsymbol{x}(0))}[\textrm{D}_{\textrm{KL}}\left(p(\cdot|\boldsymbol{x}(0))\;\|\;p_{\theta_{\text{base}}}(\cdot|\boldsymbol{x}(0))\right)].$
The KL divergence is computed according to Girsanov’s Theorem.
From here, by the maximum entropy principle applied to the marginal at time $t=1$, the first term in the objective is lower bounded by
$\textrm{D}_{\textrm{KL}}\left(p(\boldsymbol{x}(1))\;\|\;p_{\theta_{\text{base}}}(\boldsymbol{x}(1))\right)\geq\textrm{D}_{\textrm{KL}}\left(p_{\boldsymbol{\alpha}_{0}^{\ast}}(\boldsymbol{x}(1))\;\|\;p_{\theta_{\text{base}}}(\boldsymbol{x}(1))\right)$
where $p_{\boldsymbol{\alpha}_{0}^{\ast}}(\boldsymbol{x}(1))$ is the solution to the maximum entropy problem in $k$-dimensional Euclidean space. Consequently, if we can show that there exists an SDE $p(\boldsymbol{x})$ satifying $p(\boldsymbol{x}(1))=p_{\boldsymbol{\alpha}_{0}^{\ast}}(\boldsymbol{x}(1))$ and $p(\cdot|\boldsymbol{x}(1))=p_{\theta_{\text{base}}}(\cdot|\boldsymbol{x}(1))$, then $p$ is the solution to the maximum entropy problem. This is the subject of the following result:
Proposition D.2 (Maximum entropy solution for a diffusion model).
Suppose that the constraint function $\boldsymbol{h}$ depends only on the value of the path at time $t=1$ and is bounded and continuous. Moreover, assume that $\boldsymbol{x}(0)\perp\boldsymbol{x}(1)$ under the base model $p_{\theta_{\text{base}}}(\boldsymbol{x})$.
Then the solution to the maximum entropy problem is a diffusion process
$p^{\ast}:d\boldsymbol{x}(t)=\{\boldsymbol{b}_{\theta_{\text{base}}}(\boldsymbol{x}(t),t)+\sigma(t)\boldsymbol{u}^{\ast}(\boldsymbol{x}(t),t)\}dt+\sigma(t)d\boldsymbol{w}(t)$
satisfying $p^{\ast}(\boldsymbol{x}(0))=p_{\text{init}}$. The drift $\boldsymbol{u}^{\ast}(\boldsymbol{x}(t),t)$ admits the Feynman-Kac characterization
$u^{\ast}(\boldsymbol{x},t)=\sigma(t)\nabla_{\boldsymbol{x}}\log\mathbb{E}_{p_{\theta_{\text{base}}}}[\exp\{r_{\boldsymbol{\alpha}_{0}^{\ast}}(\boldsymbol{x}(1))\}|\boldsymbol{x}(t)=\boldsymbol{x}],$
where $\boldsymbol{\alpha}_{0}^{\ast}$ are the parameters corresponding to the maximum entropy solution in $k$-dimensional Euclidean space (4) with base distribution $p_{\theta_{\text{base}}}(\boldsymbol{x}(1))$ and constraint function $\boldsymbol{h}(\boldsymbol{x})$.
Finally, $p^{\ast}(\boldsymbol{x})$ satisfies $p^{\ast}(\cdot|\boldsymbol{x}(1))=p_{\theta_{\text{base}}}(\cdot|\boldsymbol{x}(1))$ and $p^{\ast}(\boldsymbol{x}(1))=p_{\boldsymbol{\alpha}_{0}^{\ast}}(\boldsymbol{x}(1))$.
We refer the reader to Domingo-Enrich et al. (2025, Theorem 1) for a proof. The result is a consequence of standard results in the theory of diffusion processes, specifically the Doob $h$-transform (Oksendal, 2013, Chapter 7).
The assumption that, under $p_{\theta_{\text{base}}}(\boldsymbol{x})$, the path at time $t=0$ is independent of the path at time $t=1$ is necessary to ensure that $p^{\ast}(\boldsymbol{x}(0))$ can be chosen to be equal to $p_{\text{init}}(\boldsymbol{x}(0))$. This is desirable because, by design, $p_{\text{init}}$ is a distribution from which sampling is tractable. However, when the independence assumption does not hold, $p^{\ast}(\boldsymbol{x}(0))$ cannot be chosen to be equal to $p_{\theta_{\text{base}}}(\boldsymbol{x}(0))$ (Denker et al., 2024, Appendix G.2).
Although, at first glance, this independence assumption may appear strong, Domingo-Enrich et al. (2025, Theorem 1) prove that diffusion models whose initial distribution is Gaussian noise and whose terminal distribution is the data distribution satisfies this property (i.e., variance-preserving SDEs). One example of a commonly used noise schedule satisfying this property is $\sigma(t)=t^{-1}$, which is singular at time $t=0$.
Section˜D.2 tells us that when we fine-tune a diffusion model with CGM to satisfy a constraint on its terminal distribution $\mathbb{E}_{p_{\theta_{\text{base}}}(\boldsymbol{x}(1))}[\boldsymbol{h}(\boldsymbol{x})]=\boldsymbol{h}^{\ast}$, we expect the terminal distribution of the base model to change to satisfy the constraint, while the conditional path distribution given the endpoint should be preserved. In other words, seeking the distribution over paths that is closest in KL distance to the base model amounts to shifting the terminal distribution while leaving the conditional distributions unchanged.
D.3 Exact Sampling from a Gaussian Mixture Model